OpenStudy (anonymous):

1. Add. (3c2 – 6c – 1) + (5c2 + 12c + 4) a. 8c2 – 6c + 3 b. 8c2 + 6c + 3 c. 8c2 – 6c – 3 d. 8c2 + 6c – 3 2. Subtract. (2f2 + 5f – 9) – (6f2 + 5) a. –4f2 – 4 b. –4f2 + 5f – 4 c. –4f2 – 14 d. –4f2 + 5f – 14

OpenStudy (anonymous):

3c2 - 6c - 1 + 5c2 + 12c + 4 3c2 + 5c2 = 8c2 -6c + 12 = 6c -1 + 4 = 3 8c2 + 6c + 3

OpenStudy (jiteshmeghwal9):

first open the brackets as \[\LARGE{3c^2-6c-1+5c^2+12c+4}\]now try to combine the like terms

OpenStudy (anonymous):

am i correct ^ ?

OpenStudy (jiteshmeghwal9):

yes ! u r correct :) Great :)

OpenStudy (anonymous):

2f2 + 5f – 9 – 6f2 + 5 2f2 - 6f2 = -4f2 + 5f -9 - 5 = -14 @jiteshmeghwal9 how about this one -4f2 + 5f - 14

hartnn (hartnn):

-9+5=-4,everything else is correct.

OpenStudy (anonymous):

okay thank you very much

hartnn (hartnn):

welcome :)